# CADS: A Systematic Literature Review on the Challenges of Abstractive Dialogue Summarization

### v2

@@ 4.1 Daily Chat
DialogSum (?) integrates dialogues from DailyDialog (?), DREAM (?), MuTual (?), and English-speaking practice websites, featuring two-speaker interactions across daily-life scenarios like work, leisure, and travel. Annotators were given guidelines on writing the summaries, such as length constraints (no longer than 20% of conversation length), maintaining an observer perspective, and explicitly describing important emotions related to events in the summary. DialogSum includes 13k dialogues, with inputs of around 1k tokens and summaries of 130 tokens.
@@ 4.2 Online Chat
SAMSum (?) comprises 16k written online dialogues from messaging apps, which linguists craft asked to create conversations similar to those they write daily, reflecting the proportion of topics of their real-life messenger conversations. The messages, each written by one person, contain chit-chats, gossip, arranging meetings, discussing politics, and consulting university assignments. They typically involve two speakers, with an average of 94 tokens per conversation, varying between three and 30 turns  [^7] A turn is the contribution made by a speaker in the form of a single utterance or a statement.. A subset of SAMSum has been adapted into formal third-person language (?) to help models transition from informal dialogue to edited text, addressing the linguistic gap between pre-training and downstream task (detailed in the language challenge, Section 3.1).
FORUM (?) consists of random samples of 100 threads from the online discussion forums on Ubuntu and TripAdvisor, with a total of 556 and 916 posts, respectively. Two human evaluators created summaries of the discussion threads, resulting in two human written summaries per sample.
CRD3 (?) consists of dialogues from the ”Critical Role Dungeon and Dragon” show with summaries collected from the Fandom wiki, featuring 159 episodes with an average dialogue length of 2550 turns.
@@ 4.3 Meeting
ICSI (?) has 59 academic group meetings with computer scientists, linguists, and engineers at the International Computer Science Institute (ICSI) in Berkeley, along with their summaries written by hired annotators. The meetings have an average of 819 turns and 13k tokens with six speakers and are about research discussions among students.
@@
ELITR (?) features 120 English technical project meetings in computer science, with each transcript averaging 7k words, 730 turns, and six speakers. The duration of the meetings varies from ten minutes to more than two hours, with an average of one hour long. Meetings shorter than half an hour are exceptions, whereas meetings longer than two hours are topic-oriented mini-workshops, also rather occasional. The original task here slightly differs as a model is required to produce not an abstractive summary but a set of meeting minutes in bullet points.
@@ 4.4 Screenplay
MediaSum (?) contains 463.6k media interview transcripts with brief summaries from National Public Radio (NPR) (?) and CNN, spanning a range of topics. The summaries are based on NPR’s interview overviews and CNN’s topic descriptions, with the latter segmented at commercial breaks to match topics to corresponding interview segments. On average, each transcript in this dataset contains 1.5k words, with summaries of around 11 words, typically involving 7 speakers.
@@ 4.5 Customer Service
TweetSumm (?) contains 1.1k dialogues derived from Twitter customer support exchanges, each with three extractive and three abstractive human-written summaries. Originating from the Kaggle Customer Support On Twitter dataset, these real-world interactions span various industries, such as airlines and retail, averaging 10 turns per dialogue and 36 tokens per abstractive summary.
@@ 4.6 Medical
MTS-Dialog (?) is a collection of 1.7k simulated doctor-patient conversations with corresponding clinical notes serving as summaries, sourced from the public Mtsamples collection (?). This dataset encompasses various medical fields, including general medicine, neurology, orthopedics, dermatology, and allergy/immunology, adhering to the SOAP (Subjective, Objective, Assessment, Plan) note format. Dialogues average 18 sentences and 242 words, with summaries typically around 81 words.
@@ 4.7 Debates
ADSC (?) features sequences of two-party dialogue chains derived from the Internet Argument Corpus (?), focusing on significant social and political topics like gun control, gay marriage, the death penalty, and abortion. It includes 45 dialogues, each accompanied by 5 unique human-generated summaries, with each summary approximately 150 words in length.
@@ 4.8 Others
FERRANTI (?) is a dataset designed for factual error correction in dialogue summarization, featuring 4k manually annotated items. The original task is to evaluate the factuality of summaries and how to correct these summaries with minimal effort. Drawing on SAMSum and DialogSum, FERRANTI includes summaries produced by models like BART (?) and UniLM (?). Human annotators assess these summaries for accuracy and identify errors, providing a focused tool for improving summary factualness.
DialSum (?) is based on the VisDial dataset (?), where two participants discuss an image from the MSCOCO dataset (?), which features approximately 120k images. Each image thereby has five captions from five different annotators.
@@ 4.9 Data Augmentation and Utility
Artificial Dataset Generation. As discussed earlier, dialogue summarization faces challenges due to the need for adaptability across various domains, structures, and speaker dynamics, which typically would be addressed by training models on diverse datasets (?). However, the scarcity and small size of datasets, with utterances sometimes containing just two to ten turns in areas like customer service and medical, restrict model training for more general applications.
Creating real-world datasets is costly and may conflict with data security and information communication policies, so artificially generated datasets are explored (?). Methods range from simple heuristic-based weak labeling, such as selecting the leading or longest utterance as proxy summary (?), to paraphrasing with updating the summary to maintain coherence (?, ?, ?), to random alterations (e.g., swapping, deletion, dialogue-act-guided insertions) and changing conversation structures (?, ?). The most common method is to use language models to create ground truth summaries either directly (?, ?, ?, ?) or after training them on human summarization techniques through few-shot learning (?).
Techniques to Maximize Dataset Utility. As training Transformer-based models requires extensive data to generalize across various topics and dialogue formats (?), researchers have explored ways to use the limited data available more effectively. A key strategy involves a prompt-based curriculum learning strategy that gradually increases the degree of prompt perturbation (e.g., word swapping, content cutting) to improve the generalization ability of models (?). Another variation is dynamic prompts, which select best-fitting few-shot samples considering dialogue content, size, and speaker number (?). Also explored are prompt transfer techniques from related dialogue domains (?) to bolster the usage of dialogue state information. Prompts are further used to split inputs into domain-invariant and domain-specific content to enhance model generalization (?, ?). Other relevant techniques include freezing most model parameters and training only specific parameters for domain adaption (?, ?, ?).
@@ 5 Evaluation
Researchers have adopted metrics from related fields and introduced new ones designed to assess the effectiveness of dialogue summarization methods. These metrics are used as a proxy for quality characteristics, such as coherence, fluency, factuality, and accuracy. In our literature review, we identify 15 metrics used more than twice in 93 papers. We categorize these metrics into four groups: count-based (e.g., ROUGE (?)), model-based (e.g., BARTScore (?)), QA-based (e.g., QAEval (?)), and human evaluation. The automatic metrics strive for alignment with human judgment. However, the correlation between these two is weak for dialogue summarization, and automatic metrics sometimes reward low-quality texts (?). Due to these recognized limitations in fully capturing the nuances of summarization quality, human evaluation is considered the gold standard across most studies. Recent analyses (?, ?) conclude that established automatic metrics may work for a superficial understanding of a model’s performance but do not align well with human judgment in discerning error nuances. These metrics either do not adequately display error occurrence or cannot reflect the impact on the quality of their scores. Hence, a composite metric of count-based, model-based, and QA-based metrics may be required for a more reliable automatic interpretation.
Section 5.1 details the characteristics typically evaluated by the metrics and Sections 5.2, 5.3, 5.4 and 5.5 further summarizes the 15 established metrics in dialogue summarization. Table 1 provides an overview of the identified metrics and their usage throughout the papers considered for this literature review.
@@ 5.1 Assessed Characteristics
Metrics assess the quality of a generated text, often in relation to a reference text. Low-quality text may contain more redundancy, incoherence, grammatical errors, poor structure, or inappropriate language, while high-quality text would closely align with the reference. We identify ten key characteristics (in italic) organized into four overarching groups: accuracy, content, readability, and style.
Accuracy is the core of a high-quality summary (?), making adherence to the original text’s truth and, therefore, factuality paramount. A summary must accurately reflect the facts, events, and details from the source without distortion.
@@
Readability is essential for making summaries accessible and easy to comprehend. Coherence ensures that information is presented logically with smooth transitions, making the summary easy to follow (?). The structure and organization of a summary further enhances its readability and facilitates information processing, ensuring that it is well-organized and logically constructed (?). Conciseness (?) and non-redundancy (?) are also essential as a summary should contain only the essence of the input transcript without unnecessary details.
Style further contributes to the perceived quality of a text. Consistency in perspective, tense, and style throughout the summary contributes to its readability and professional quality (?). Fluency focuses on grammar, vocabulary, and sentence structure and aims for a natural, easy-to-read summary that is free from phrasing errors, enhancing the information’s overall clarity and accessibility (?).
@@ 5.2 Count-Based
Count-based metrics, including N-gram-based measures like BLEU (?), and ROUGE (?), and statistical measures such as Perplexity (?), are often static and rule-based algorithms that have a long history for evaluating text summarization. However, due to their simple approach, they have been criticized for their limitations in capturing overall meaning, fluency, coherence, or factuality in their scores (?).
BLEU (?) measures the precision of generated summaries against reference texts by examining the overlap of word sequences (typically 1- to 3-word N-grams (?)), emphasizing similarity to the original phrasing but risking a linear relationship to noise (?).
@@
METEOR (?) advances BLEU and ROUGE by incorporating both precision and recall, along with synonym matching for semantic analysis, thus enabling to capture the semantic similarity between a candidate and a reference. Despite its more balanced approach and sentence-level focus, METEOR, like its predecessors, is prone to noise interference (?).
@@ 5.3 Model-Based
With advancements in language models, there is an increasing focus on model-based evaluation metrics due to their higher correlation with human judgment, though count-based metrics remain popular. Model-based metrics range from those using handcrafted features and learned embeddings to hybrids combining contextual embeddings with count-based approaches. After representing the respective text in a latent space, these metrics consider the similarity between the candidate and reference text, aiming to offer a more nuanced assessment by focusing on semantic similarity, likelihood, and usefulness of summaries. However, despite their potential to adapt to changes in language use (e.g., distribution drifts) (?), these metrics can still be error-prone (?), slower than count-based metrics and may not directly measure the specific characteristics outlined in Section 5.1, making it challenging to discern which particular aspect influences their scores.
BERTScore (?) leverages BERT embeddings (?) to contextually embed the reference and candidate text, calculate the pairwise cosine-similarity on the token level to build a similarity matrix, and greedily pick the highest similarity scores to compute the harmonic mean of precision and recall, thereby capturing semantic nuances (?).
MoverScore (?) was introduced concurrently to BERTScore following a similar approach, as it also leverages BERT embeddings (?) but instead considers the distance between reference and candidate text, hence employing the Word Mover’s Distance (?), a special case of the Earth Mover’s Distance (?), to measure semantic distance. This change in measure allows MoverScore to map semantically related words from one sequence to one word in the other sequence (many-to-one).
BARTScore (?) evaluates the plausibility of generating a reference text from a given generated text, and vice versa, by calculating the log-likelihood of a sequence that a BART (?) model would typically generate based on the given context. This evaluation focuses on assessing both the fluency and the semantic accuracy of the generated content.
@@ 5.4 QA-Based
The QA-based metrics we identify in the literature concentrate on the factual correctness of a summary using external question-answering systems. The effectiveness of these metrics largely depends on the performance of the QA systems, which may not always align with human evaluation standards.
@@
QuestEval (?) combines the approaches of FEQA and SummQA, thereby adopting a bidirectional strategy, generating questions from the summary to compare with the source text and vice versa. This bidirectional evaluation offers a balanced and holistic assessment by considering included and omitted information in the summary.
@@
Count-based N-gram BLEU word overlap, precision, multiple references 21
ROUGE word overlap, recall, one reference 127
METEOR word overlap, precision and recall, one reference 3
@@
[…]
Model-based Hybrid BERTScore token similarity, cosine-similarity, one-to-one 38
MoverScore token similarity, mover distance, one-to-many 3
FactCC entailment classifier, scoring based on consistency 9
[…]
BARTScore mimics human judgement, promptable 13
BLANC fill-in-the-blanc task with and without the summary 2
QA-based FEQA questions based on summary, answered through input 38
@@
QuestEval combination of FEQA and SummQA 3
Human Evaluation Performance Likert Scale scale, typically 1 (worst) to 5 (best) 6
@@
[…]
Cohen’s Kappa measure agreements, categorical data, 2 raters 3
Fleiss’ Kappa measure agreement, nominal data, 2 raters 2
Table 1: Relevant metrics and evaluation measures employed in dialogue summarization with more than twice reported use. ’Usage’ states the number of papers reporting the evaluation measure in their publication.

### v3

@@ 4.1 Daily Chat
DialogSum (?) integrates dialogues from DailyDialog (?), Dream (?), Mutual (?), and English-speaking practice websites, featuring two-speaker interactions across daily-life scenarios like work, leisure, and travel. Annotators were given guidelines on writing summaries, such as length constraints (no longer than 20% of conversation length). DialogSum includes 13k dialogues, with inputs of around 1k tokens and summaries of 130 tokens.
@@ 4.2 Online Chat
SAMSum (?) comprises 16k written online dialogues from messaging apps, which linguists craft asked to create conversations similar to those they write daily, reflecting the proportion of topics of their real-life messenger conversations. The messages, each written by one person, contain chit-chats, gossip, arranging meetings, discussing politics, and consulting university assignments. They typically involve two speakers, with an average of 94 tokens per conversation, varying between three and 30 turns [^8] A turn is a contribution made by a speaker in the form of a single utterance or a statement.. A subset of SAMSum has been adapted into formal third-person language (?) to help models transition from informal dialogue to edited text, addressing the linguistic gap between pre-training and downstream task (detailed in the language challenge, Section 3.1).
Forum (?) consists of random samples of 100 threads from the online discussion forums on Ubuntu and TripAdvisor, with a total of 556 and 916 posts, respectively. Two human evaluators created summaries of the discussion threads, resulting in two human-written summaries per sample.
Crd3 (?) consists of dialogues from the ‘Critical Role Dungeon and Dragon’ show with summaries collected from the Fandom wiki, featuring 159 episodes with an average dialogue length of 2550 turns.
@@ 4.3 Meeting
ICSI (?) has 59 academic group meetings with computer scientists, linguists, and engineers at the International Computer Science Institute (ICSI) in Berkeley, along with their summaries written by hired annotators. The meetings have an average of 819 turns and 13k tokens with six speakers and are research discussions among students.
@@
Elitr (?) features 120 English technical project meetings in computer science, with each transcript averaging 7k words, 730 turns, and six speakers. The duration of the meetings varies from ten minutes to more than two hours, with an average of one hour long. Meetings shorter than half an hour are exceptions, whereas meetings longer than two hours are topic-oriented mini-workshops, also rather occasional. The original task of the dataset differs from abstractive summarization as a model is required to produce not an abstractive summary but a set of meeting minutes in bullet points.
@@ 4.4 Screenplay
MediaSum (?) contains 463.6k media interview transcripts with summaries from National Public Radio (NPR, ?, 2020) and CNN, spanning a range of topics, including politics, news, crime, and economy. The summaries are based on NPR’s interview overviews and CNN’s topic descriptions, with the latter segmented at commercial breaks to match topics to corresponding interview segments. On average, each transcript in this dataset contains 1.5k words, with summaries of around 11 words, typically involving seven speakers.
@@ 4.5 Customer Service
TweetSumm (?) contains 1.1k dialogues derived from Twitter customer support exchanges, each with three extractive and three abstractive human-written summaries. Originating from the Kaggle Customer Support On Twitter dataset, these real-world interactions span various industries, such as airlines and retail, averaging ten turns per dialogue and 36 tokens per abstractive summary.
@@ 4.6 Medical
MTS-Dialog (?) is a collection of 1.7k simulated doctor-patient conversations with corresponding clinical notes serving as summaries, sourced from the public Mtsamples collection (?). This dataset encompasses various medical fields, including general medicine, neurology, orthopedics, dermatology, and immunology, adhering to the SOAP (Subjective, Objective, Assessment, Plan) note format. Dialogues average 18 sentences and 242 words, with summaries typically around 81 words.
@@ 4.7 Debates
ADSC (?) features sequences of two-party dialogue chains derived from the Internet Argument Corpus (?), focusing on significant social and political topics like gun control, gay marriage, the death penalty, and abortion. It includes 45 dialogues, each accompanied by five unique human-generated summaries, with each summary approximately 150 words in length.
@@ 4.8 Others
Ferranti (?) is a dataset designed for factual error correction in dialogue summarization, featuring 4k manually annotated items. The original task is to evaluate the factuality of summaries and how to correct these summaries with minimal effort. Drawing on SAMSum and DialogSum, Ferranti includes summaries produced by models like BART (?) and Unilm (?). Annotators assess these summaries for accuracy and identify errors, providing a focused tool for improving summary factualness.
DialSum (?) is based on the VisDial dataset (?), where two participants discuss images from the MSCOCO dataset (?), which features $\sim$120k images. Each image has five captions from five different annotators.
Type Dataset Challenge Descriptive Tags Usage
Daily Chat DialogSum Speaker daily life scenarios 26
Online SAMSum Speaker, Salience messaging apps 68
Chat FORUM Salience threads 5
CRD3 Language live-streamed show 6
Meeting AMI Salience, Comprehension staged business meetings 63
ICSI Salience, Language academic group meetings 21
QMSum Salience, Language query-based summarization 18
Elitr Salience technical project meetings in computer science 7
MeetingBank Structure parliament meetings 7
Screenplay MediaSum Structure media interview 15
SummScreen Speaker, Salience TV series transcripts 8
Customer TweetSumm Language Twitter customer support 7
Service TODSum Language, Factuality open-domain task-oriented dialogues 6
Medical MTS-Dialog Comprehension simulated doctor-patient conversations with clinical notes 5
Debates ADSC Comprehension two-party dialogues on social and political topics 5
Others ConvoSumm Salience dialogues from comments, emails, and threads 5
Ferranti Factuality human assessed automatic summaries 5
DialSum Comprehension two-party discussion on images 10
Table 3: Matching between established datasets and their primary reported challenge.
‘Usage’ indicates the number of papers reporting the dataset in their publication.
@@ 4.9 Data Augmentation and Utility
Artificial Dataset Generation. As discussed earlier, dialogue summarization faces challenges due to the need for adaptability across various domains, structures, and speaker dynamics, which typically would be addressed by training models on diverse datasets (?). However, datasets’ scarcity and small size, with utterances containing just two to ten turns in areas like customer service and medical, restrict model training for more general applications.
Creating real-world datasets is costly and may conflict with data security and information communication policies, so artificially generated datasets are explored (?). Methods range from simple heuristic-based weak labeling, such as selecting the leading or longest utterance as proxy summary (?), to paraphrasing with updating the summary to maintain coherence (?, ?, ?), to random alterations (e.g., swapping, deletion, dialogue-act-guided insertions) and changing conversation structures (?, ?). The most common method is to use language models to create ground truth summaries either directly (?, ?, ?, ?) or after training them on human summarization patterns through few-shot learning (?).
Techniques to Maximize Dataset Utility. As the training of Transformer-based models requires extensive data to generalize across various topics and dialogue formats (?), researchers have explored ways to use the limited data available more effectively. A key strategy involves a prompt-based curriculum learning strategy that gradually increases the degree of prompt perturbation (e.g., word swapping, content cutting) to improve the generalization ability of models (?). Another variation are dynamic prompts, which select best-fitting few-shot samples considering dialogue content, size, and speaker number (?). Also explored are prompt transfer techniques from related dialogue domains (?) to bolster the usage of dialogue state information (i.e., data used to represent the underlying intentions and goals within a dialogue). Prompts are further used to split inputs into domain-invariant and domain-specific content to enhance model generalization (?, ?). Further techniques include freezing most model parameters and training only specific parameters for domain adaption (?, ?, ?).
@@ 5 Evaluation
Researchers have adopted metrics from related fields and introduced new ones designed to assess the effectiveness of dialogue summarization methods. These metrics proxy for quality characteristics, such as coherence, fluency, factuality, and accuracy. Our literature review identifies 15 metrics used more than twice in 93 papers. We categorize these metrics into four groups: count-based (e.g., ROUGE ?, 2004), model-based (e.g., BARTScore ?, 2021), QA-based (e.g., QAEval ?, 2021), and human evaluation. The automatic metrics strive for alignment with human judgment. However, the correlation between these two is weak for dialogue summarization, and automatic metrics sometimes reward low-quality texts (?). Due to these recognized limitations in fully capturing the nuances of summarization quality, human evaluation is considered the gold standard across most studies. Recent analyses (?, ?) conclude that established automatic metrics may work for a superficial understanding of a model’s performance but do not align well with human judgment in discerning error nuances. These metrics show individual weaknesses where they individually do not adequately reflect occurrence across all error types (e.g., of hallucinated content) or cannot show the impact on the quality in their scores (e.g., when information is omitted). Hence, a composite metric of count-based, model-based, and QA-based metrics may be required for a more reliable automatic interpretation, with the individual metrics focusing on specific error types.
Section 5.1 details the characteristics typically evaluated by metrics and Sections 5.2, 5.3, 5.4 and 5.5 further summarizes the 15 established metrics in dialogue summarization. Table 4 provides an overview of the identified metrics and their usage throughout the papers considered for this literature review.
@@ 5.1 Assessed Characteristics
Metrics assess the quality of a generated text, often in relation to a reference text. Low-quality text may contain more redundancy, incoherence, grammatical errors, poor structure, or inappropriate language, while high-quality text would closely align with the reference. We identify ten key characteristics discussed in the literature (in italic) and organize them into four overarching groups: accuracy, content, readability, and style.
Accuracy is the core of a high-quality summary (?), making it crucial to maintain the truth of the original text and ensure factuality. A summary must accurately reflect the facts, events, and details from the source without any distortion.
@@
Readability is essential for making summaries accessible and easy to grasp the key points quickly. Coherence ensures that information is presented logically with smooth transitions, making the summary easy to follow (?). The structure and organization of a summary further enhances its readability and facilitates information processing, ensuring that it is well-organized and logical (?). Conciseness (?) and non-redundancy (?) are also essential as a summary should contain only the essence of the input transcript without unnecessary details.
Style pertains to the text’s perceived quality and formal presentation, contributing to its professional polish. Consistency in perspective, tense, and stylistic choices throughout the summary contributes to the overall perception of the text’s professionalism (?). Fluency focuses on grammar, vocabulary, and sentence structure and aims for a natural, easy-to-read summary free from phrasing errors, enhancing the information’s overall clarity and accessibility (?).
@@ 5.2 Count-Based
Count-based metrics, including N-gram-based measures like BLEU (?), and ROUGE (?), and statistical measures such as Perplexity (?), are often static and rule-based algorithms that have a long history for evaluating text summarization. However, due to their simplicity, they have been criticized for their limitations in capturing overall meaning, fluency, coherence, or factuality (?).
BLEU (?) measures the precision of generated summaries against reference texts by examining the overlap of word sequences (typically 1- to 3-word N-grams, ?, 2019), emphasizing similarity to the original phrasing but risking a linear relationship to noise (?).
@@
METEOR (?) advances BLEU and ROUGE by incorporating both precision and recall, along with synonym matching for semantic analysis, thus enabling the capturing of the semantic similarity between a candidate and a reference. Despite its more balanced approach and sentence-level focus, METEOR, like its predecessors, is prone to noise interference (?).
@@ 5.3 Model-Based
With advancements in language models, there is an increasing focus on model-based evaluation metrics due to their higher correlation with human judgment, though count-based metrics remain popular. Model-based metrics encompass a wide range of approaches, including those that calculate semantic similarity (e.g., BERTScore, MoverScore), text generation likelihood (e.g., BARTScore), and entailment probability (e.g., FactCC). These metrics typically represent text in a latent space using pre-trained embeddings or contextual representations, aiming to provide a more nuanced assessment by focusing on semantic similarity, likelihood of text generation, and factual consistency. While these metrics have the potential to adapt to evolving language use (e.g., distribution drifts) (?), they can still be error-prone (?), slower than count-based metrics, and may not directly measure specific characteristics (as outlined in Section 5.1), making it challenging to discern which particular aspect influences their scores.
BERTScore (?) leverages BERT embeddings (?) to assess textual similarity. The metric first contextually embeds the reference and candidate texts, then constructs a similarity matrix through pairwise cosine similarities on the token level. The final score is computed by greedily selecting the highest similarity scores and calculating the harmonic mean of precision and recall, enabling the metric to capture semantic nuances beyond surface-level matching (?).
MoverScore (?) was introduced concurrently to BERTScore following a similar approach, as it also leverages BERT embeddings but instead considers the distance between reference and candidate text, hence employing the Word Mover’s Distanc (?), a special case of the Earth Mover’s Distance (?), to measure semantic distance. This change in measure allows MoverScore to map semantically related words from one sequence to one word in another sequence (many-to-one).
BARTScore (?) evaluates the plausibility of generating a reference text from a given generated text, and vice versa, by calculating the log-likelihood of a sequence that a BART (?) model would typically generate based on the given context. This evaluation focuses on assessing both the fluency and the semantic accuracy.
@@ 5.4 QA-Based
The QA-based metrics we identify in the literature focus on the factual correctness of a summary by using external question-answering systems. These metrics leverage pre-trained transformer-based models to generate questions and assess whether the summary contains the correct answers. Their effectiveness largely depends on the quality of the underlying QA systems, which may not always align perfectly with human evaluation standards.
@@
QuestEval (?) combines the approaches of FEQA and SummaQA, thereby adopting a bidirectional strategy, generating questions from the summary to compare with the source text and vice versa. This bidirectional evaluation offers a balanced and holistic assessment by considering included and omitted information in the summary.
@@
Count based N-gram Bleu overlap, precision, multiple references 21
Rouge overlap, recall, one reference 127
Meteor overlap, precision and recall, one reference 3
@@
Statistical Perplexity likelihood of word sequences 3
Model based Hybrid Bertscore token-level, cosine similarity, one-to-one 38
Moverscore token-level, mover distance, one-to-many 3
Factcc entailment classifier, scores consistency 9
Trained Bleurt mimics human judgment 7
Bartscore mimics human judgment, promptable 13
Blanc fill-in-the-blank task with and without the summary 2
QA based Feqa questions based on summary, answered through input 38
@@
Questeval combination of Feqa and Summaqa 3
Human Evaluation Performance Likert Scale ordinal scale, e.g., 1 (worst) to 5 (best) 6
@@
Agreement Krippendorff’s Alpha disagreement, different data formats 2
Cohen’s Kappa agreements, categorical data, 2 raters 3
Fleiss’ Kappa agreement, nominal data, 2 raters 2
Table 4: Relevant metrics and evaluation measures employed in dialogue summarization with more than twice reported use. ‘Usage’ states the number of papers reporting the evaluation measure in their publication.